OpenStudy (anonymous):

Find what amount 125 is 25% of using the decimal approach.

OpenStudy (anonymous):

simple rule (just write it right, and use cross-rule): 125 - 25% x - 100% x = (125*100)/25 = 500

OpenStudy (anonymous):

125 = 0.25x x = (125)/(0.25) = 500

OpenStudy (anonymous):

thank you. :)